Eddie has 12 feet of fencing to enclose a rectangular garden. If the garden’s length is 10 feet less than 3 times its width, then the area of the garden is:

Answer by Alan3354(69443)

Eddie has 12 feet of fencing to enclose a rectangular garden. If the garden’s length is 10 feet less than 3 times its width, then the area of the garden is:
--------------
L = 3W-10
2L + 2W = 12
L + W = 6
(3W-10) + W = 6
4W = 16
W = 4
L = 2
Area = 8 sq feet